A Bibliography of Publications of Raj Jain

Raj Jain The Ohio State University

Dept of Computer and Information Science 2015 Neil Avenue, DL395 Columbus, OH 43210-1277

USA

Tel: +1 614 292 3989 FAX: +1 614 292 2911

E-mail: jain@acm.org (Internet)

13 October 2017 Version 1.09

Abstract

This bibliography records publications of Raj Jain.

Title word cross-reference

P^{2} [JC85].

18th [Wil82].

21st [IEE98].

’88 [ACM88].

’97 [IEE97a].

ABR [FJK^{+}96, FJK^{+}98, FJG^{+}98, Jai95a, Jai95c, JKF^{+}96, KJFG96b, KJF^{+}96a, KJFG96a, KJG^{+}96, KJF^{+}96b, KJJ^{+}97, KJF^{+}97b, KJF^{+}97a, KJGF97, Kal97, KJJ^{+}98, KVJ^{+}98, KJG^{+}98, KJF^{+}00, VKJ^{+}97, VJGF99]. abstract [ZJ86]. access [AJSI10]. Accounting [JT82]. Acknowledgments [GJK^{+}97c]. Adapter [Jai91c]. Address [Jai90a, Jai89b, Jai92a, Jai92b, Jaixx]. Advances [Jai96]. Ahead [Jai95b]. Algorithm [JC85, KJF^{+}97b, KJF^{+}00]. Algorithms [CJ89, FJG^{+}98, Jai86a]. All-Optical [ABB^{+}93]. Allocation [JCH84]. Analysis [CJ89, JH86, Jai90d, Jai90e, Jai91a, Jai91b, Jai91c, VJGF99]. annual [IEE98]. App [Ano95]. Applications [DJ96]. Approach [Jai78, Jai89a]. April [IEE88, IEE98]. Architectural [KGJ97]. architecture [SIJPP11, JH86]. area [IEE88]. Art [Jai91a, Jai91b]. Asynchronous [Kal97]. ATM [FJK^{+}96, FJK^{+}97, FJK^{+}98, FJG^{+}98, GJK^{+}97b, GJK^{+}97c, GJK^{+}97a, GJF^{+}98, Jai95b, Jai95a, Jai96, JKV96, JKF^{+}96, JB97, JKVxx, KJF^{+}96a, KJFG96a, KJG^{+}96, KJJ^{+}97, KJF^{+}97b, KJF^{+}97a, KJGF97, Kal97, KJJ^{+}98, KVJ^{+}98, KJG^{+}98, KJF^{+}00, KGJ97, SJ95, VJGF99]. ATM-UBR [GJK^{+}97b, GJK^{+}97c, GJK^{+}97a]. Automation [Jai74]. Available [Kal97, KJG^{+}98]. AVC [AJSI10]. Avoidance [CJ89, JRC87, JR88, Jai89a, JKV96, JKVxx, RCJ87, RJ88, RJ90, RJ95].

backbone [Ano95, KJF^{+}97a]. Background [KVJ^{+}98]. Band [ABB^{+}93]. Bandwidth [FJK^{+}98]. Based [Jai86b, Jai89a]. Behavior [JKF^{+}96]. Binary [RCJ87, RJ88, RJ90, RJ95]. Bit [GJF^{+}98, Kal97, KJG^{+}98]. Boston [Wea78]. Brief [SJ95]. broadband [AJSI10]. Buffer [KJFG96b]. Buffering [KJF^{+}96a]. Bursty [VKJ^{+}97].

CA [ACM88, IEE98]. Caching [Jai85, Jai90a]. Calculation [JC85]. Canada [IEE97b, IEE97a]. century [IEE98]. Challenges [Jai95b]. Characteristics [Jai90a, Jai90c]. Characterization [JT82]. Communications [FJK^{+}97, IEE95, IEE97a, M^{+}97, IEE96, IEE98]. Comparison [Jai90a, Jai89b, Jai92a, Jai92b, Jaixx, ZJ86]. Computer [CJ89, IEE88, Jai78, JCH84, JR86, JRC87, JR88, Jai90a, Jai89b, Jai89a, Jai90b, Jai91a, Jai91b, Jai92a, Jai92b, Jaixx, M^{+}97, RCJ87, RJ88, RJ90, Wea78, Wil82, IEE98, RJ95]. concept [SIJPP11]. Concepts [JR88, HJ04]. Conference [IEE95, IEE97b, IEE97a, M^{+}97, IEE98]. Congestion [CCJ95, CJ89, FJLK95, Jai86b, JRC87, JR88, Jai89a, Jai90b, Jai92c, Jai96, JKV96, JKVxx, RCJ87, RJ88, RJ90, LJ01, RJ95]. Congrès [IEE97b]. conjunction [IEE97b]. Connectionless [JRC87, JR88, RCJ87, RJ88, RJ95]. Connections [FJK^{+}98, FJG^{+}98].

Considerations [KJJ^{+}97, KGJ97, KJJ^{+}98]. Consolidation [FJG^{+}98]. Consortium [ABB^{+}93]. Control [CCJ95, FJLK95, Jai78, Jai79, Jai86b, Jai90b, Jai96, VJGF99]. Control-theoretic [Jai78, Jai79]. Controlled [Jai85, Jai86b]. Converter [Jai74]. CPEUG [Wea78, Wil82]. Current [Jai93].

Data [FJLK95, Jai90c, Jai95c]. DC

[IEE88, Wil82]. Delay [Jai89a]. Dependent [KVJ^{+}98]. Design [GJF^{+}98, Jai91b, KJJ^{+}97, KJJ^{+}98, VJGF99]. Destination [Jai90a, KJJ^{+}97, KJJ^{+}98]. Determining [FJK^{+}98]. Digital [JH86]. Discrimination [JCH84]. Distributed [Jai90c]. Divergence [Jai86a]. Drop [GJK^{+}97c]. Dynamic [JC85, VJGF99].

Effect [Jai90d, Jai90e]. Effort [JB97]. Emulation [SJ95]. ENM [IEE97b]. ENM-97 [IEE97b]. Enterprise [IEE97b]. ERICA [KJF^{+}97b, KJF^{+}00]. Error [Jai90c]. Evaluation [Wea78, Wil82]. Experimental [Jai91b]. Explanation [JKF^{+}96]. Explicit [CCJ95, JKVxx, LJ01]. Extensions [JKV96].

Fair [FJK^{+}98]. Fairness [JCH84]. FDDI [Jai90c, Jai90d, Jai90e, Jai91c, Jai93, Jai94b, Jai94a]. Feature [KJJ^{+}97, KJJ^{+}98]. Feedback [FJG^{+}98, RCJ87, RJ88, RJ90, RJ95]. Fiber [Jai90c, Jai94b, Jai94a]. First [IEE97b]. Flies [Jai95c]. Flow [Jai85, Jai86b]. Flow- [Jai86b]. Formulation [Jai79]. Forum [JB97]. fourteenth [Wea78]. Francisco [IEE98]. front [LJ01]. functions [VJGF99]. Future [Jai93, SIJPP11]. FWNs [SIJPP11].

gateway [IEE98]. General [RCJ87].

generation [AJSI10]. Global

[DJ96, IEE96]. GLOBECOM [IEE96]. Goals [JR88]. GPS [DJ96]. Group [Wea78, Wil82]. Guarantees [GJF^{+}98]. Guidelines [Jai90d, Jai90e].

Handbook [Jai94b, Jai94a]. Hashing [Jai89b, Jai92a, Jai92b, Jaixx]. held [Wea78]. Heterogeneous [Jai89a]. High [HJ04, Jai91c, Jai92c, Jai94b, Ano95, Jai94a]. High-Speed [Jai91c, Jai94b, Ano95, Jai94a]. Histograms [JC85]. Hotel [IEE88, IEE98].

ICC [IEE97b, IEE97a]. ICC-97 [IEE97b]. ICC’95 [IEE95]. ID [SIJPP11]. ID/ Locator [SIJPP11]. Image [JT82]. Improve [GJK^{+}97c]. Improvement [Jai78]. Improving [GJK^{+}97b, GJK^{+}97a, LJ01]. Increase [CJ89]. Increase/Decrease [CJ89]. Indication [CCJ95, JKVxx]. INFOCOM’98 [IEE98]. Interconnected [Jai89a]. Interface [Jai90c]. International [IEE95, IEE97a, M^{+}97]. Internet [KJF^{+}96a]. Issues [FJK^{+}97, GJF^{+}98, Jai90b, Jai93, Jai95b, HJ04]. IV [RCJ87].

joint [IEE98]. June [IEE97b, IEE97a].

key [IEE96]. Killer [Ano95]. knowledge [IEE97a].

LAN [SJ95]. Layer [JRC87, JR88, RCJ87, RJ88, RJ95]. Layers [SJ95]. LD [Jai74]. Learnt [JKV96]. Lessons [JKV96]. Link [FJLK95]. Links [FJK^{+}96]. LISP [ZJ86]. Locality [Jai90a]. London [IEE96]. Long [KVJ^{+}98]. Long-Range [KVJ^{+}98]. Lookup [Jai89b, Jai92a, Jai92b, Jaixx]. Lose [KJGF97, KJG^{+}98]. Lose-it [KJG^{+}98].

machines [ZJ86]. Management [Jai79, Jai92c, Jai96, JKF^{+}96, KJF^{+}97b, Kal97, SJ95, KJF^{+}00]. March [IEE98]. mark [LJ01]. mark-front [LJ01]. Massachusetts [Wea78]. Measure [JCH84]. Measurement [Jai91b]. Measurements [JR86]. Media [Jai94b, Jai94a]. meeting [Wea78, Wil82]. Methodology [JR88]. millennium [IEE97a]. Mini [IEE97b]. Mini-Conference [IEE97b]. Minimum [GJF^{+}98]. mobile [AJSI10]. Mode [Kal97]. Model [JR86]. Modeling [AJSI10, JH86, Jai91b]. Models [Jai74]. Montréal [IEE97b, IEE97a]. Multipoint [FJK^{+}97, FJG^{+}98]. Myths [Jai92c].

National [IEE88]. Network [Jai85, JR86, JRC87, JR88, KGJ97, RCJ87, RJ88, RJ95]. Networking [DJ96, IEE88, IEE97b, Jai94b, HJ04, Jai94a, JH86]. Networks [ABB^{+}93, CJ89, FJLK95, FJK^{+}96, FJK^{+}98, FJG^{+}98, GJK^{+}97c, Jai86b, JRC87, JR88, Jai90a, Jai89b, Jai89a, Jai90b, Jai90d, Jai90e, Jai92a, Jai92b, Jai92c, Jai95b, Jai95a, Jai96, JKV96, Jaixx, KJFG96a, KJG^{+}96, KJJ^{+}97, KJF^{+}97b, KJGF97, Kal97, KJG^{+}98, M^{+}97, RCJ87, RJ88, RJ90, SIJPP11, AJSI10, JKVxx, KJJ^{+}98, KVJ^{+}98, KJF^{+}00, RJ95]. Nevada [M^{+}97]. Nikko [IEE98]. notification [LJ01]. November [IEE96].

Observations [JC85]. October [Wea78, Wil82]. Open [FJK^{+}97]. Operating [Jai79]. Optical [ABB^{+}93]. Order [Jai85]. OSU [JKV96, JKVxx]. Other [Jai94b, Jai94a]. Out-of-Order [Jai85]. Overview [JB97, SJ95].

Packet [Jai86a, JR86]. Packets [Jai85]. Palais [IEE97b]. Parameters [Jai90d, Jai90e]. Part [RCJ87]. patterns [KJF^{+}97a]. Decrease [CJ89]. IP [HJ04, KJFG96b, KJFG96a, KJG^{+}96, KJF^{+}96b, KGJ97]. Locator [SIJPP11]. VD [KJJ^{+}97, KJJ^{+}98]. Virtual [KJJ^{+}97, KJJ^{+}98]. Performance [GJK^{+}97b, GJK^{+}97c, GJK^{+}97a, Jai78, JH86, Jai90d, Jai90e, Jai91a, Jai91b, Jai91c, JB97, KJF^{+}96a, KJFG96a, KJG^{+}96, KJF^{+}96b, KJF^{+}97a, KVJ^{+}98, KGJ97, VKJ^{+}97, Wea78, Wil82, HJ04, ZJ86]. Point [FJG^{+}98]. Point-to-Multipoint [FJG^{+}98]. Policies [GJK^{+}97c, Jai79, KJGF97, KJG^{+}98]. Positioning [DJ96]. Potential [DJ96]. Precompetitive [ABB^{+}93]. Proceedings [IEE98, Wil82, IEE88, M^{+}97, Wea78]. prosperity [IEE96]. Protocol [SJ95, VC94]. Protocols [FJLK95, FJK^{+}97, KJF^{+}96a]. providing [GJF^{+}98].

Quantiles [JC85]. Quantitative [JCH84]. Québec [IEE97b, IEE97a]. queue [VJGF99].

Range [KVJ^{+}98]. Rate [CCJ95, GJF^{+}98, JKVxx, Kal97, KJG^{+}98]. record [IEE97a]. Requirements [KJFG96b, KJF^{+}96a]. Resource [Jai79, JCH84]. Retransmissions [Jai86a]. Ring [Jai90d, Jai90e]. Rules [FJK^{+}96].

San [IEE98]. Satellite [FJLK95, FJK^{+}96, GJK^{+}97c, KJG^{+}96, KVJ^{+}98, KGJ97]. Scheme [Jai86b, JKV96, JKVxx, RCJ87, RJ88, RJ90, RJ95]. Schemes [Jai90a, Jai89b, Jai92a, Jai92b, Jaixx, VJGF99]. Seattle [IEE95]. Selective [GJK^{+}97c, RCJ87]. September [M^{+}97]. Service [FJK^{+}96, GJF^{+}98, Jai95a, KJFG96a, KJJ^{+}97, KJGF97, Kal97, KJG^{+}98, GJK^{+}97b, GJK^{+}97a, Jai95c, KJJ^{+}98]. Services [KJF^{+}96a, KJG^{+}96]. Setting [Jai90d, Jai90e]. seventeenth [IEE98]. Share [FJK^{+}98]. Shared [JCH84]. Sheraton [IEE88]. SIGCOMM [ACM88]. Simulation [Jai91b]. Sixth [M^{+}97]. Societies [IEE98]. solutions [HJ04]. Source [FJK^{+}96, JKF^{+}96, KJJ^{+}97, KJJ^{+}98]. Source/Virtual [KJJ^{+}97, KJJ^{+}98].

Sources [VKJ^{+}97]. specification [VC94]. Speed [Jai91c, Jai92c, Ano95, Jai94b, Jai94a]. split [SIJPP11]. Standby [Jai95c]. Stanford [ACM88]. Storing [JC85]. strategy [LJ01]. Survey [FJLK95, FJK^{+}97, Jai96, KJGF97]. SVC [AJSI10]. SVC-TS [AJSI10]. Switch [KJF^{+}97b, KJF^{+}00, VJGF99]. Symposium [ACM88, IEE88]. Systems [DJ96, Jai78, Jai79, JCH84, Jai91a, Jai91b, ZJ86].

TCP [GJK^{+}97b, GJK^{+}97c, GJK^{+}97a, HJ04, KJFG96b, KJFG96a, KJG^{+}96, KJF^{+}96b, KJF^{+}97a, KVJ^{+}98, KGJ97]. TCP/IP [HJ04, KJFG96b, KJFG96a, KJG^{+}96, KJF^{+}96b, KGJ97]. Techniques [FJLK95, Jai91b]. Terrestrial [GJK^{+}97c, KVJ^{+}98]. Testing [JB97, VC94]. theoretic [Jai78, Jai79]. Timeout [Jai86a, Jai86b]. Token [Jai90d, Jai90e]. Topologies [RCJ87]. traces [AJSI10]. Traffic [JR86, Jai96, JKF^{+}96, KJF^{+}97b, Kal97, KVJ^{+}98, SJ95, KJF^{+}97a, KJF^{+}00]. Trains [JR86]. Trains-Measurements [JR86]. Transfer [Kal97]. Trends [Jai90b, Jai93]. TS [AJSI10]. TTRT [Jai90d, Jai90e]. two [ZJ86].

UBR [GJK^{+}97b, GJK^{+}97c, GJK^{+}97a, KJF^{+}96a, KJG^{+}96]. Unspecified [GJF^{+}98]. USA [IEE95, IEE98]. Use [KJGF97, KJG^{+}98]. Use-it [KJG^{+}98]. Use-It-Or-Lose-It [KJGF97]. Users [Wea78, Wil82]. Using [JT82, JKVxx, KJG^{+}96, Jai94b, Jai94a, SIJPP11].

various [KJF^{+}97a]. VBR

[KJF^{+}97a, KVJ^{+}98]. Vegas [M^{+}97]. verification [VC94]. video [AJSI10]. Virtual [KJJ^{+}97, KJJ^{+}98]. Virtualization [SIJPP11]. VS [KJJ^{+}97, KJJ^{+}98]. VS/VD [KJJ^{+}97, KJJ^{+}98].

WA [IEE95]. Washington [IEE88, Wil82].

Web [VKJ^{+}97]. Wide [ABB^{+}93, VKJ^{+}97]. Wide-Band [ABB^{+}93]. Window [Jai85, Jai86b, ZJ86]. Wireless [SIJPP11]. Without [JC85]. Workload [JT82]. World [VKJ^{+}97]. WWW [VKJ^{+}97].

XIV [VC94].

References

Alexander:1993:PCW

[ABB^{+}93] S. B. Alexander, R. S. Bondurant, D. Byrne, V. W. S. Chan, S. G. Finn, et al. A precompetitive consortium on wideband all-optical networks. Journal of lightwave technology, 11(5– 6):714–735, May/June 1993. CO-DEN JLTEDG. ISSN 0733-8724 (print), 1558-2213 (electronic).

URL  http://www.cis.ohio-
state.edu/~jain/papers/aon.

zip. Special issue on broad-band optical networks.

ACM:1988:ASS

[ACM88] ACM, editor. ACM SIGCOMM ’88 Symposium (1988: Stanford, CA), volume 18(4) of Computer communications review. ACM Press, New York, NY 10036, USA, August 1988. ISBN 0-89791-279-9. LCCN TK 5105.5 C655 v.18 no.4 1988.

AlTamimi:2010:MGA

[AJSI10] Abdel Karim Al Tamimi, Raj Jain, and Chakchai So-In. Modeling and generation of AVC and SVC-TS mobile video traces for broadband access networks. In MMSys ’10: Proceedings of the first annual ACM SIGMM conference on Multimedia systems, February 2010, pages 89– 98. ACM Press, New York, NY 10036, USA, 2010. ISBN 1-60558-914-4. LCCN ????

Anonymous:1995:KAH

[Ano95] Anonymous. Killer app: It’s the high-speed backbone. Government computer news, ?? (??):24, October 16, 1995. ISSN 0738-4300. URL http:

//www.cis.ohio-state.edu/~
jain/papers/gcn.htm. An inter-

view with Raj Jain.

Charny:1995:CCE

[CCJ95] A. Charny, D. D. Clark, and R. Jain. Congestion control with explicit rate indication. In IEEE [IEE95], pages 1954–1963. ISBN 0-7803-2486-2, 0-7803-2487-0, 0-7803-2488-9. ISSN 0536-1486. LCCN TK5101.A1 I244 1995. URL

http://www.cis.ohio-state.
edu/~jain/papers/charny.htm.

Three volumes.

Chiu:1989:AID

[CJ89] Dah-Ming Chiu and Raj Jain. Analysis of the increase/decrease algorithms for congestion avoidance in computer networks. Computer Networks and ISDN Systems, 17(1):1–14, June 1989. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Dommety:1996:PNA

[DJ96] G. Dommety and Raj Jain. Potential networking applications of global positioning systems (GPS). Technical Report TR-24, ????, ????, April 1996. ??

pp. URL http://www.cis.ohio-
state.edu/~jain/papers/gps.
htm.

Fahmy:1998:FCA

[FJG^{+}98] Sonia Fahmy, Raj Jain, Rohit Goyal, Bobby Vandalore, Shivkumar Kalyanaraman, Sastri Kota, and Pradeep Samudra. Feedback consolidation algorithms for ABR point-to-multipoint connections in ATM networks. In IEEE [IEE98], pages 1004–1013. ISBN 0-7803-4383-2 (softbound), 0-7803-4384-0 (casebound), 0-7803-4385-9 (microfiche). LCCN TK5105.5 .I33 1998. URL http:

//www.cis.ohio-state.edu/~
jain/papers/cnsldt.htm. Three

volumes.

Fahmy:1996:SRA

[FJK^{+}96] Sonia Fahmy, Raj Jain, Shivkumar Kalyanaraman, Rohit Goyal, and Fang Lu. On source rules for ABR service on ATM networks with satellite links. In Proceedings of the First International Workshop on Satellite-based Information Services, Rye, New York, November 1996, pages 108–115. ????, ????, 1996. URL http:

//www.cis.ohio-state.edu/~
jain/papers/wosbis.htm.

Fahmy:1997:SPO

[FJK^{+}97] Sonia Fahmy, Raj Jain, Shivkumar Kalyanaraman, Rohit Goyal, Bobby Vandalore, and Xiangrong Cai. A survey of protocols and open issues in ATM multipoint communications. Technical Report 47, The Ohio State University, Department of Computer and Information Science, Columbus, OH 43210-1277, USA, August 21, 1997. ?? pp. URL http:

                                        //www.cis.ohio-state.edu/~
                                        jain/papers/mcast.htm.

Fahmy:1998:DFB

[FJK^{+}98] Sonia Fahmy, Raj Jain, Shivkumar Kalyanaraman, Rohit Goyal, and Bobby Vandalore. On determining the fair bandwidth share for ABR connections in ATM networks. In Proceedings of the IEEE International Conference on Communications (ICC), page 7. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 1998. URL http:

                                        //www.cis.ohio-state.edu/~
                                        jain/papers/neff.htm.

Fahmy:1995:SCC

[FJLK95] Sonia Fahmy, Raj Jain, Fang Lu, and Shivkumar Kalyanaraman. A survey of congestion control techniques and data link protocols in satellite networks. International Journal of Satellite Communications, ?? (??):31, 1995. CODEN IJSCEF. ISSN 0737-2884. URL http:

                                        //www.cis.ohio-state.edu/~
                                        jain/papers/sat_surv.htm.

Goyal:1998:DIP

[GJF^{+}98] Rohit Goyal, Raj Jain, Sonia Fahmy, Bobby Vandalore, and Shivkumar Kalyanaraman. Design issues for providing minimum rate guarantees to the ATM unspecified bit rate service. In IEEE [IEE98], pages 169–175. ISBN 0-7803-4383-2 (softbound), 0-7803-4384-0 (casebound), 0-7803-4385-9 (microfiche). LCCN TK5105.5 .I33 1998. URL http:

//www.cis.ohio-state.edu/~
jain/papers/gfr.htm. Three

volumes.

Goyal:1997:UIP

[GJK^{+}97a] Rohit Goyal, Raj Jain, Shiv Kalyanaraman, Sonia Fahmy, and Seong-Cheol Kim. UBR+: Improving performance of TCP over ATM-UBR service. In IEEE [IEE97b], pages 1042–1048. ISBN 0-7803-4112-0 (softbound), 0-7803-4113-9 (microfiche). LCCN TK5105.7.I3233 1997. URL http:

//www.cis.ohio-state.edu/~
jain/papers/icc97.htm.

Goyal:1997:IPT

[GJK^{+}97b] Rohit Goyal, Raj Jain, Shiv Kalyanaraman, Sonia Fahmy, and Bobby Vandalore. Improving the performance of TCP over the ATM-UBR service. Computer Communications, ??(??):??, July 1997. CODEN COCOD7. ISSN 0140-3664 (print), 1873-703X (electronic). URL http:

//www.cis.ohio-state.edu/~
jain/papers/cc.htm.

Goyal:1997:TSA

[GJK^{+}97c] Rohit Goyal, Raj Jain, Shivkumar Kalyanaraman, Sonia Fahmy, Bobby Vandalore, and Sastri Kota. TCP selective acknowledgments and UBR drop policies to improve ATM-UBR performance over terrestrial and satellite networks. In Makki et al. [M^{+}97], pages 17–27. ISBN 0-8186-8186-1, 0-8186-8188-8 (microfiche). LCCN TK5105.5 .I5712 1997. URL http:

                                        //www.cis.ohio-state.edu/~
                                        jain/papers/ic3n97.htm.

Hassan:2004:HPT

[HJ04] Mahbub Hassan and Raj Jain, editors. High performance TCP/IP networking: concepts, issues, and solutions. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-064634-2. xx + 383 pp. LCCN TK5105.585 .H54 2003. US$59.00.

IEEE:1988:CNS

[IEE88] IEEE, editor. Computer Networking Symposium: proceedings, Sheraton National Hotel, Washington, DC area, April 11– 13, 1988. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0835-8 (paperback), 0-8186-8835-1 (case), 0-8186-4835-X (microfiche). LCCN TK5105.5 .C6441 1988. IEEE catalog number 88CH2547-8. Computer Society order number 835.

IEEE:1995:III

[IEE95] IEEE, editor. ICC’95: IEEE International Conference on Communications (1995 : Seattle, WA, USA). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2486-2, 0-7803-2487-0, 0-7803-2488-9. ISSN 0536-1486. LCCN TK5101.A1 I244 1995. Three volumes.

IEEE:1996:IGL

[IEE96] IEEE, editor. IEEE GLOBECOM 1996: London, November 18–22, 1996: communications: the key to global prosperity. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3336-5 (softbound), 0-7803-3337-3 (casebound), 0-7803-3338-1 (microfiche). LCCN TK 5101 A1 I146 1996a.

IEEE:1997:IIC

[IEE97a] IEEE, editor. 1997 IEEE International Conference on Communications: towards the knowledge millennium: ICC ’97, 8–12 June 1997, Montréal, Québec, Canada, conference record. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-3925-8 (softbound), 0-7803-3926-6 (casebound), 0-7803-3927-4 (microfiche), 0-7803-3928-2 (CD-ROM). LCCN TK5101.A1 I16 1997. Three volumes.

IEEE:1997:FIE

[IEE97b] IEEE, editor. First IEEE Enterprise Networking Mini-Conference (ENM-97) in conjunction with the ICC-97, June 11–12, 1997, Palais des Congrès, Montréal, Québec, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-4112-0 (softbound), 0-7803-4113-9 (microfiche). LCCN TK5105.7.I3233 1997.

IEEE:1998:PII

[IEE98] IEEE, editor. Proceedings: IEEE INFOCOM’98: the conference on computer communications: seventeenth annual joint conference of the IEEE Computer and Communications Societies: gateway to the 21st century, 29 March– 2 April 1998, Hotel Nikko, San Francisco, CA, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-7803-4383-2 (softbound), 0-7803-4384-0 (casebound), 0-7803-4385-9 (microfiche). LCCN TK5105.5 .I33 1998. Three volumes.

Jain:1974:MLC

[Jai74] R. Jain. Models of LD converter for its automation. In Proc. Symposium on Industrial Automation, Bangalore, India, page ?? ????, ????, June 1974.

Jain:1978:CTA

[Jai78] R. Jain. Control-theoretic approach to computer systems performance improvement. In Weatherbee [Wea78], pages 93– 100. LCCN QC100 .U57 no. 500-41; QA76.9.E94.

Jain:1979:CTF

[Jai79]  Rajendra K. Jain.  Control-

theoretic Formulation of Operating Systems Resource Management Policies. Outstanding Dissertations in the Computer Sciences Series. Garland Publishing Company, New York, NY, USA, 1979. ISBN 0-8240-4402-9. 215 pp. LCCN QA76.9.E94J34 1980. Originally presented as the author’s thesis, Harvard University, 1978.

Jain:1985:COP

[Jai85] R. Jain. On caching out-of-order packets in window flow controlled network. DEC Research Report DEC-TR-342, Eastern Research Laboratory, Digital Equipment Corporation, 77 Reed Road, Hudson, MA 01749, USA, September 1985. 8 pp.

Jain:1986:DTA

[Jai86a] R. Jain. Divergence of timeout algorithms for packet retransmissions. In Proceedings of the 5th IEEE Phoenix Conference on Computers and Communication, pages 174–179. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, March 1986.

Jain:1986:TBC

[Jai86b] R. Jain. A timeout based congestion control scheme for window flow- controlled networks. IEEE Journal on Selected Areas in Communications, SAC-4(7): 1162–1167, October 1986. CO-DEN ISACEM. ISSN 0733-8716 (print), 1558-0008 (electronic). Reprinted in C. Partridge, Ed., Innovations in Internetworking, 289–295, Artech House, Norwood, MA 1988.

Jain:1989:DBA

[Jai89a] R. Jain. A delay based approach for congestion avoidance in interconnected heterogeneous computer networks. Computer Communications Review, ACM SIGCOMM, ??(??):56–71, October 1989. CODEN CCRED2. ISSN 0146-4833. URL http:

                                        //www.cis.ohio-state.edu/~
                                        jain/papers/delay.htm.

Jain:1989:CHS

[Jai89b] Raj Jain. A comparison of hashing schemes for address lookup in computer networks. Technical Report DEC-TR-593, Digital Equipment Corporation, February 1989. ?? pp.

Jain:1989:CDA

[Jai90a] R. Jain. Characteristics of destination address locality in computer networks: A comparison of caching schemes. Computer Networks and ISDN Systems, 18: 243–254, 1989/1990. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

                                        http://www.cis.ohio-state.
                                        edu/~jain/papers/cache_ps.
                                        htm.

Jain:1990:CCC

[Jai90b] R. Jain. Congestion control in computer networks: Trends and issues. IEEE network, pages 24–30, May 1990. ISSN 0890-8044. URL http:

                                        //www.cis.ohio-state.edu/~
                                        jain/papers/cong_trends.htm.

Jain:1990:ECF

[Jai90c] R. Jain. Error characteristics of fiber distributed data interface (FDDI). IEEE Transactions on Communications, 38(8): 1224–1252, August 1990. CO-DEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

URL  http://www.cis.ohio-
state.edu/~jain/papers/xie1.
htm.

Jain:1990:PAFa

[Jai90d] R. Jain. Performance analysis of FDDI token ring networks: Effect of parameters and guidelines for setting TTRT. IEEE Lightwave Telecommunication Systems, 20 (2):16–22, May 1990. CODEN IELTEU. ISSN 1055-6877. URL

http://www.cis.ohio-state.
edu/~jain/papers/ttrt_lcs_

1col.htm. Also in Proceedings ACM SIGCOMM’90, Philadelphia, PA, September 24–27, 1990, pp. 264–275 (Computer communication review, Vol. 20, No. 4).

Jain:1990:PAFb

[Jai90e] R. Jain. Performance analysis of FDDI token ring networks: Effect of parameters and guidelines for setting TTRT. Computer Communications Review, ACM SIG-COMM, 20(4):264–275, September 1990. CODEN CCRED2. ISBN 0-89791-405-8. ISSN 0146-4833.

Jain:1991:ACSa

[Jai91a]  R. Jain.  The art of com-

puter systems performance analysis. Computer Communications, 14(4):253–??, May 1991. CO-DEN COCOD7. ISSN 0140-3664 (print), 1873-703X (electronic).

Jain:1991:ACSb

[Jai91b] R. Jain. The Art of Computer Systems Performance Analysis: Techniques for Experimental Design, Measurement, Simulation, and Modeling. Wiley-Interscience, New York, NY, USA, May 1991. ISBN 0-471-50336-3. xxvii + 685 pp. LCCN QA76.9.E94J32 1991. Winner of “1991 Best Advanced How-To Book, Systems” award from the Computer Press Association.

Jain:1991:PAH

[Jai91c] Raj Jain. Performance analysis of a high-speed FDDI adapter. Digital Technical Journal, 3(3): 1–11 (??), Summer 1991. CO-DEN DTJOEL. ISSN 0898-901X.

                                        URL ftp://ftp.digital.com/
                                        pub/Digital/info/DTJ/v3n3/
                                        Performance_Analysis_of_FDDIgh_
                                        01apr1992DTJ307P8.ps.

Jain:1992:CHSa

[Jai92a] R. Jain. A comparison of hashing schemes for address lookup in computer networks. Technical report, DEC, 1992. 17 pp. URL

                                        http://www.cis.ohio-state.
                                        edu/~jain/papers/hash_ps.htm.

Jain:1992:CHSb

[Jai92b] R. Jain. A comparison of hashing schemes for address lookup in computer networks. IEEE Transactions on Communications, 40(3):1570–1573, October 1992. CODEN IECMBT.

ISSN 0090-6778 (print), 1558-0857 (electronic). URL http:

//www.cis.ohio-state.edu/~
jain/papers/hash_ieee_2col.
htm.

Jain:1992:MAC

[Jai92c] R. Jain. Myths about congestion management in high speed networks. Internetworking: Research and Experience, 3(??):101–113, ???? 1992. CODEN IREEE7. ISSN 1049-8915. URL http:

//www.cis.ohio-state.edu/~
jain/papers/cong_myths.htm.

Jain:1993:FCI

[Jai93] R. Jain. FDDI: Current issues and future trends. IEEE Communications Magazine, 31(9):98– 105, September 1993. CODEN ICOMD9. ISSN 0163-6804. URL

http://www.cis.ohio-state.
edu/~jain/papers/gmu_paper.
htm.

Jain:1994:FHH

[Jai94a] R. Jain. FDDI handbook: highspeed networking using fiber and other media. Addison-Wesley, Reading, MA, USA, April 1994. ISBN 0-201-56376-2. xxvii + 528 pp. LCCN TK5105.72 .J35 1994.

Jain:1994:FH

[Jai94b] Raj Jain. FDDI Handbook: Highspeed Networking using Fiber and Other Media. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-56376-2. xxvii + 528 pp. LCCN TK5105.72 .J35 199. US$49.50.

Jain:1995:ASA

[Jai95a] R. Jain. ABR service on ATM networks: What is it? Network World, ??(??):??, June 24, 1995. ISSN 0887-7661 (print), 1944-7655 (electronic). URL http:

//www.cis.ohio-state.edu/~
jain/papers/networkw.htm.

Jain:1995:ANI

[Jai95b] R. Jain. ATM networks: Issues and challenges ahead. In Vuong and Chanson [VC94], pages 12–?? ISBN 0-412-63640-9. LCCN TK 5105.55 P76 1994. URL http:

//www.cis.ohio-state.edu/~
jain/papers/atm_issu.htm. Pa-

pers from the PSTV ’94 Symposium, organized under the auspices of IFIP W.G. 6.1 and held in Vancouver, BC, Canada, in August 1994.

Jain:1995:DFS

[Jai95c] R. Jain. Data flies standby with ABR service. Network World, ??(??):43, June 12, 1995. ISSN 0887-7661 (print), 1944-7655 (electronic). URL http:

//www.cis.ohio-state.edu/~
jain/papers/networkw.htm.

Jain:1996:CCT

[Jai96] Raj Jain. Congestion control and traffic management in ATM networks: Recent advances and A survey. Computer Networks and ISDN Systems, 28(13):1723– 1738, October 15, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.cis.ohio-state.
edu/~jain/papers/cnis.htm;

http://www.elsevier.com/cgi-
bin/cas/tree/store/comnet/
cas_sub/browse/browse.cgi?
year=1996&volume=28&issue=
13&aid=1565.

Jain:19xx:CHS

[Jaixx] Raj Jain. A comparison of hashing schemes for address lookup in computer networks. Technical Report DEC-TR-566, Digital Equipment Corporation, 19xx. ?? pp.

URL jain%erlang.dec@decwrl.
dec.com.

Jain:1997:PTE

[JB97] R. Jain and G. Babic. Performance testing effort at the ATM forum: An overview. IEEE Communications Magazine, 35(8): 110–116, August 1997. CODEN ICOMD9. ISSN 0163-6804. URL

http://www.cis.ohio-state.
edu/~jain/papers/perf_com.
htm.

Jain:1985:ADC

[JC85] Raj Jain and Imrich Chlamtac.

The P^{2} algorithm for dynamic calculation of quantiles and histograms without storing observations. Communications of the ACM, 28(10):1076–1085, October 1985. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/
0001-0782/4378.html.

Jain:1984:QMF

[JCH84] R. Jain, D. Chiu, and W. Hawe.

A quantitative measure of fairness and discrimination for resource allocation in shared computer systems. DEC Research Report TR-301, Digital Equipment Corporation, Maynard, MA, USA, September 1984. 38 pp.

URL ftp://ftp.netlab.ohio-
state.edu/pub/jain/papers/
fairness.htm.

Jain:1986:PAM

[JH86] R. Jain and W. R. Hawe. Performance analysis and modeling of Digital’s Networking Architecture. Digital Technical Journal, 1 (3):25–34, September 1986. CO-DEN DTJOEL. ISSN 0898-901X.

Jain:1996:SBA

[JKF^{+}96] R. Jain, S. Kalyanaraman, S. Fahmy, R. Goyal, and S. Kim. Source behavior for ATM ABR traffic management: An explanation. IEEE Communications Magazine, 34(11):50–57, November 1, 1996. CODEN ICOMD9. ISSN 0163-6804. URL http:

//www.cis.ohio-state.edu/~
jain/papers/src_rule.htm.

Jain:1996:OSC

[JKV96] Raj Jain, Shiv Kalyanaraman, and Ram Viswanathan. The OSU scheme for congestion avoidance in ATM networks: Lessons learnt and extensions. Performance Evaluation, 31(1–2):67–88, May 1996. CODEN PEREDN. ISSN 0166-5316 (print), 1872-745X (electronic). URL http:

//www.cis.ohio-state.edu/~
jain/papers/perf-sub.htm.

Jain:19xx:OSC

[JKVxx] Raj Jain, Shiv Kalyanaraman, and Ram Viswanathan. The OSU scheme for congestion avoidance in ATM networks using explicit rate indication. In Proceedings WATM’95 First Workshop on ATM Traffic Management, Paris, France, page 20. ????, ????, 19xx. URL http:

//www.cis.ohio-state.edu/~
jain/papers/osuc_con.htm.

Jain:1986:PTM

[JR86] R. Jain and S. Routhier. Packet trains-measurements and a new model for computer network traffic. IEEE Journal on Selected Areas in Communications, SAC-4(6):986–995, September 1986. CODEN ISACEM. ISSN 0733-8716 (print), 1558-0008 (electronic). URL http:

//www.cis.ohio-state.edu/~
jain/papers/train.htm. Reprinted

in Amit Bhargava, Ed., Integrated Broadband Networks, Artech House, Norwood, MA, 1990.

Jain:1988:CAC

[JR88] R. Jain and K. Ramakrishnan.

Congestion avoidance in computer networks with A connectionless network layer: Concepts, goals, and methodology. In IEEE [IEE88], pages 134– 143. ISBN 0-8186-0835-8 (paperback), 0-8186-8835-1 (case), 0-8186-4835-X (microfiche). LCCN TK5105.5 .C6441 1988. IEEE catalog number 88CH2547-8. Computer Society order number 835.

Jain:1987:CAC

[JRC87] R. Jain, K. Ramakrishnan, and D. Chiu. Congestion avoidance in computer networks with a connectionless network layer. Technical Report DEC-TR-506, DEC, 1987. 17 pp. URL http:

                                        //www.cis.ohio-state.edu/~
                                        jain/papers/cr5.htm. DEC-

TR-506, reprinted in C. Partridge, Ed., Innovations in Internetworking, 140–156, published by Artech House, October 1988.

Jain:1982:WCU

[JT82] R. Jain and R. Turner. Workload characterization using image accounting. In Wilson [Wil82], pages 111–120. LCCN QC100 .U57 no. 500-95; QA76.9.E94.

Kalyanaraman:1997:TMA

[Kal97] Shivkumar Kalyanaraman. Traffic Management for the Available Bit Rate (ABR) Service in Asynchronous Transfer Mode (ATM) Networks. PhD thesis, The Ohio State University, Columbus, OH, USA, 1997. xxiv + 429 pp. URL

                                        http://www.cis.ohio-state.
                                        edu/~jain/theses/shiv.htm.

Kota:1997:SAN

[KGJ97] Sastri Kota, Rohit Goyal, and Raj Jain. Satellite ATM network architectural considerations and TCP/IP performance. In Proceedings of the 3rd Ka Band Utilization Conference, Sorrento, Italy, 1997, pages 481–488. Istituto Internazionale delle Comunicazioni, Genova, Italy, 1997. URL http:

                                        //www.cis.ohio-state.edu/~
                                        jain/papers/kaband.htm.

Kalyanaraman:1996:PBR

[KJF^{+}96a] Shiv Kalyanaraman, Raj Jain, Sonia Fahmy, Rohit Goyal, and Seong-Cheol Kim. Performance and buffering requirements of internet protocols over ATM ABR and UBR services. IEEE Communications Magazine, 36(6):152– 157, September 1996. CODEN ICOMD9. ISSN 0163-6804. URL

http://www.cis.ohio-state.
edu/~jain/papers/ieee-mag.
htm.

Kalyanaraman:1996:PTIb

[KJF^{+}96b] Shiv Kalyanaraman, Raj Jain, Sonia Fahmy, Rohit Goyal, Fang Lu, and Saragur Srinidhi. Performance of TCP/IP over ABR. In IEEE [IEE96], pages 468–475. ISBN 0-7803-3336-5 (softbound), 0-7803-3337-3 (casebound), 0-7803-3338-1 (microfiche). LCCN TK 5101 A1 I146 1996a. URL

http://www.cis.ohio-state.
edu/~jain/papers/globecom.
htm.

Kalyanaraman:1997:PTA

[KJF^{+}97a] Shiv Kalyanaraman, Raj Jain, Sonia Fahmy, Rohit Goyal, Jianping Jiang, and Seong-Cheol Kim. Performance of TCP over ABR on ATM backbone and with various VBR traffic patterns. In IEEE [IEE97a], pages 1035–1041. ISBN 0-7803-3925-8 (softbound), 0-7803-3926-6 (casebound), 0-7803-3927-4 (microfiche), 0-7803-3928-2 (CD-ROM). LCCN TK5101.A1 I16 1997. URL http:

//www.cis.ohio-state.edu/~

jain/papers/tcp_vbr.htm. Three

volumes.

Kalyanaraman:1997:ESA

[KJF^{+}97b] Shivkumar Kalyanaraman, Raj Jain, Sonia Fahmy, Rohit Goyal, and Bobby Vandalore. The ER-ICA switch algorithm for ABR traffic management in ATM networks. Technical report, The Ohio State University, Department of Computer and Information Science, Columbus, OH 43210-1277, USA, November 1997. 32 pp. URL http:

//www.cis.ohio-state.edu/~
jain/papers/erica.htm.

Kalyanaraman:2000:ESA

[KJF^{+}00] Shivkumar Kalyanaraman, Raj Jain, Sonia Fahmy, Rohit Goyal, and Bobby Vandalore. The ERICA switch algorithm for ABR traffic management in ATM networks. IEEE/ACM Transactions on Networking, 8(1): 87–98, February 2000. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (elec-

tronic). URL http://www.acm.
org/pubs/citations/journals/
ton/2000-8-1/p87-kalyanaraman/

.

Kalyanaraman:1996:PTI

[KJFG96a] S. Kalyanaraman, R. Jain, S. Fahmy, and R. Goyal. Performance of TCP/IP over ABR service on ATM networks. In IEEE [IEE96], pages 468–475. ISBN 0-7803-3336-5 (softbound), 0-7803-3337-3 (casebound), 0-7803-3338-1 (microfiche). LCCN TK 5101 A1 I146 1996a.

Kalyanaraman:1996:BRT

[KJFG96b] Shiv Kalyanaraman, Raj Jain, Sonia Fahmy, and Rohit Goyal. Buffer requirements for TCP/ IP over ABR. In Proc. IEEE ATM’96 Workshop, San Francisco, August 23–24, 1996, page 6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. URL http:

//www.cis.ohio-state.edu/~
jain/papers/atm96.htm.

Kalyanaraman:1996:PTIa

[KJG^{+}96] Shiv Kalyanaraman, Raj Jain, Rohit Goyal, Sonia Fahmy, and Seong-Cheol Kim. Performance of TCP/IP using ATM ABR and UBR services over satellite networks. In Proc. IEEE Communication Society Workshop on Computer-Aided Modeling , Analysis and Design of Communication Links and Networks, McLean, VA, October 20, 1996, page 8. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 1996. URL http:

//www.cis.ohio-state.edu/~
jain/papers/satellit.htm.

Kalyanaraman:1998:UIL

[KJG^{+}98] Shivkumar Kalyanaraman, Raj Jain, Rohit Goyal, Sonia Fahmy, and Pradeep Samudra. Use-it or lose-it policies for the available bit rate (ABR) service in ATM networks. Computer Networks and ISDN Systems, 30(24):2293–2308, December 14, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

                                        http://www.cis.ohio-state.
                                        edu/~jain/papers/uili.htm;
                                         http://www.elsevier.com/cas/
                                        tree/store/comnet/sub/1998/
                                        30/24/1964.pdf.

Kalyanaraman:1997:SUI

[KJGF97] Shivkumar Kalyanaraman, Raj Jain, Rohit Goyal, and Sonia Fahmy. A survey of the use-it-or-lose-it policies for the ABR service in ATM networks. Technical Report OSU-CISRC-1/97-TR02, Ohio State University, Columbus, OH, USA, May 1997. ??

                                        pp. URL http://www.cis.ohio-
                                        state.edu/~jain/papers/tr_
                                        rule5.htm.

Kalyanaraman:1997:DCV

[KJJ^{+}97] Shiv Kalyanaraman, Raj Jain, Jianping Jiang, Rohit Goyal, Sonia Fahmy, and Pradeep Samudra. Design considerations for the virtual source/virtual destination (VS/VD) feature in the ABR service of ATM networks. In IEEE [IEE98], pages 21–?? ISBN 0-7803-4383-2 (softbound), 0-7803-4384-0 (casebound), 0-7803-4385-9 (microfiche). LCCN TK5105.5 .I33 1998. URL http:

                                        //www.cis.ohio-state.edu/~
                                        jain/papers/vsvd.htm. Three

volumes.

Kalyanaraman:1998:DCV

[KJJ^{+}98] Shivkumar Kalyanaraman, Raj Jain, Jianping Jiang, Rohit Goyal, and Sonia Fahmy. Design considerations for the virtual source/virtual destination (VS/ VD) feature in the ABR service of ATM networks. Computer Networks and ISDN Systems, 30(19):1811–1824, October 14, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

URL  http://www.elsevier.
com/cas/tree/store/comnet/
sub/1998/30/19/1970.pdf.

Kalyanaraman:1998:PTA

[KVJ^{+}98] Shivkumar Kalyanaraman, Bobby Vandalore, Raj Jain, Rohit Goyal, Sonia Fahmy, Seong-Cheol Kim, and Sastri Kota. Performance of TCP over ABR with longrange dependent VBR background traffic over terrestrial and satellite ATM networks. In IEEE [IEE98], page ?? ISBN 0-7803-4383-2 (softbound), 0-7803-4384-0 (casebound), 0-7803-4385-9 (microfiche). LCCN TK5105.5 .I33 1998. URL http:

//www.cis.ohio-state.edu/~
jain/papers/mpeg.htm. Three

volumes.

Liu:2001:IEC

[LJ01] Chunlei Liu and Raj Jain. Improving explicit congestion notification with the mark-front strategy. Computer Networks (Amsterdam, Netherlands: 1999), 35(2–3):185–201, February ??, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/10/
15/22/52/27/28/abstract.html;
 http://www.elsevier.nl/gej-
ng/10/15/22/52/27/28/article.
pdf.

Makki:1997:SIC

[M^{+}97] Kia Makki et al., editors. Sixth International Conference on Computer Communications and Networks, September 22–25, 1997, Las Vegas, Nevada: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-8186-1, 0-8186-8188-8 (microfiche). LCCN TK5105.5 .I5712 1997.

Ramakrishnan:1987:CAC

[RCJ87] K. K. Ramakrishnan, D. Chiu, and R. Jain. Congestion avoidance in computer networks with a connectionless network layer, part IV: A selective binary feedback scheme for general topologies. Technical Report DEC-TR-510, DEC, August 1987. 43 pp.

                                        URL ftp://ftp.netlab.ohio-
                                        state.edu/pub/jain/papers/
                                        dectr510.pdf.

Ramakrishnan:1988:BFS

[RJ88] K. Ramakrishnan and R. Jain. A binary feedback scheme for congestion avoidance in computer networks with connectionless network layer. In ACM [ACM88], pages 303–313. ISBN 0-89791-279-9. LCCN TK 5105.5 C655 v.18 no.4 1988. Selected as the best paper.

Ramakrishnan:1990:BFS

[RJ90] K. K. Ramakrishnan and R. Jain. A binary feedback scheme for congestion avoidance in computer networks. ACM Transactions on Computer Systems, 8 (2):158–181, May 1990. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1990-8-2/p158-
ramakrishnan/.

Ramakrishnan:1995:BFS

[RJ95] K. K. Ramakrishnan and Raj Jain. A binary feedback scheme for congestion avoidance in computer networks with a connectionless network layer. SIG-COMM Computer Communication Review, 25(1):??, January 1995.

So-In:2011:VAU

[SIJPP11] Chakchai So-In, Raj Jain, Subharthi Paul, and Jianli Pan. Virtualization architecture using the ID/Locator split concept for Future Wireless Networks (FWNs). Computer Networks (Amsterdam, Netherlands: 1999), 55(2):415– 430, February 1, 2011. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Siu:1995:BOA

[SJ95] K. Siu and R. Jain. A brief overview of ATM: Protocol layers, LAN emulation, and traffic management. Computer Communications Review, ACM SIGCOMM, 25(2):6–28, April 1995. CO-DEN CCRED2. ISSN 0146-4833.

URL  http://www.cis.ohio-
state.edu/~jain/papers/atm_
tut.htm.

Vuong:1994:PST

[VC94] Son T. Vuong and Samuel T. Chanson, editors. Protocol specification, testing and verification XIV. Chapman and Hall, New York, NY, USA, 1994. ISBN 0-412-63640-9. LCCN TK 5105.55 P76 1994. Papers from the PSTV ’94 Symposium, organized under the auspices of IFIP W.G. 6.1 and held in Vancouver, BC, Canada, in August 1994.

Vandalore:1999:DQC

[VJGF99] Bobby Vandalore, Raj Jain, Rohit Goyal, and Sonia Fahmy. Dynamic queue control functions for ATM ABR switch schemes: design and analysis. Computer Networks (Amsterdam, Netherlands: 1999), 31(18):1935– 1949, August 17, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL  http://www.elsevier.
                                        com/cas/tree/store/comnet/
                                        sub/1999/31/18/2133.pdf.

Vandalore:1997:PBW

[VKJ^{+}97] Bobby Vandalore, Shivkumar Kalyanaraman, Raj Jain, Rohit Goyal, Sonia Fahmy, and Seong-Cheol Kim. Performance of bursty World Wide Web (WWW) sources over ABR. In WebNet ‘97, Toronto, November 1997, page ?? ????, ????, 1997. URL http:

                                        //www.cis.ohio-state.edu/~
                                        jain/papers/webspec.htm.

Weatherbee:1978:CPE

[Wea78] James E. Weatherbee, editor. Computer Performance Evaluation Users Group (CPEUG): proceedings of the fourteenth meeting held at Boston, Massachusetts, October 24–27, 1978, volume 500-41 of NBS special publication. U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1978. LCCN QC100 .U57 no. 500-41; QA76.9.E94.

Wilson:1982:PCP

[Wil82] Carol B. Wilson, editor. Proceedings of the Computer Performance Evaluation Users Group (CPEUG) 18th meeting, Washington, DC, October 25–28, 1982, volume 500-95 of NBS special publication. U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1982. LCCN QC100 .U57 no. 500-95; QA76.9.E94.

Zabih:1986:PCW

[ZJ86] Ramin Zabih and Raj Jain. A performance comparison of the window systems of two LISP machines (abstract). In CSC ’86: Proceedings of the 1986 ACM fourteenth annual conference on Computer science, February 1986, page 458. ACM Press, New York, NY 10036, USA, 1986. ISBN 0-89791-177-6. LCCN ????